OpenStudy (anonymous):

You have red, blue, and green marbles in the ratio of 5:4:3, and you have a total of 36 marbles. How many blue marbles do you have?

OpenStudy (studygurl14):

5:4:3 5 + 4 + 3 = 12 total blue marbles: 4 4/12 \[\frac{ 4 }{ 12 } = \frac{ B }{ 36 }\]

OpenStudy (anonymous):

4

OpenStudy (studygurl14):

4(36) = 12B Can you solve the rest?